Operation Summary for July 7, 2010 - July 9, 2010
Wednesday July 7
The day shift began with Operations monitoring the Tevatron (TeV) store 7956 (124.57E30) and the Recycler stash (208.3E10), with the Antiproton Source (Pbar) stacking (11.1E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Operators prepared T-977 for access during the morning for alignment work - while this was going on, one-second slow spill beam studies were conducted. After the alignment work was finished, T-977 resumed taking beam. Operations terminated store 7956 at 1:04 PM. Operations put the TeV into shot setup for a 3x3 (protons by antiprotons) studies store, but the store was cut due to problems with a Linac RF (LRF4) station. Operations put the TeV through a dry squeeze. Investigation and repair of LRF4 lasted through to the end of the shift. FESS personnel repaired the leaking AC unit in the Linac gallery while there was no beam.
During the evening shift, Linac resumed sending beam at 5:10 PM. Operations put the TeV into shot setup for another 3x3 studies store at 5:30 PM, and established store 7958 at 6:21 PM with an average initial luminosity (AIL) of 11.7E30. Operations terminated the studies store at 9:56 PM, and put the TeV into HEP shot setup. The setup lasted through to the end of the shift due to several small problems with the Tevatron Electron Lens (TEL1).
Thursday July 8
The midnight shift began with the TeV in shot setup, Operations monitoring the stash (3.9E10), with Pbar stacking (42.5E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established store 7959 at 12:10 AM with an AIL of 261.3E30. Operations established beam to T-977 at 4 AM. Operations reported some minor downtime due to various Linac RF trips. At 4:30 AM, the TeV sector C49 vertical separator sparked and caused a 4.5E30 loss in luminosity.
During the day shift, MI and MINERvA personnel conducted a neutrino communications study, which limited beam to NuMI and Pbar between 10:10 AM and 12:33 PM. Operations reported some minor downtime due to various Linac RF trips.
During the evening shift, Operations disabled the T-977 beam at 6:13 PM. Operations terminated store 7959 at 8:10 PM, put the TeV into shot setup, and established store 7960 at 9:35 PM with an AIL of 325.5E30. Operations reported some downtime when the Booster west gallery glycol water skid was inadvertently tripped off, taking down the anode supply. Operators recovered the station, but BRF18 would not reset; operators bypassed the station. Operators accessed enclosures C, D, & E in Switchyard (SY) to check on sump alarms.
Friday July 9
The midnight shift began with Operations monitoring store 7960 (194.59E30) and the stash (84.0E10), with Pbar stacking (13.1E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations reported that a power supply (S:HP3US) would not turn on after the SY access – the supply would trip off with an overcurrent fault and a loud bang. Operators called in EE Support personnel to make repairs; they had the supply up and running by 2:22 AM. Operations established beam to T-977 at 4:16 AM. The Pbar Debuncher bend bus current dropped by ~15 amps or so, which halted stacking. Pbar personnel investigated the problem.
The Plans for Friday and the Weekend
The plans for today are to get Pbar back to stacking and for the weekend are to stack, stash, and store. NuMI will begin their six-week shutdown on Monday morning (7/12/10) at 4 AM.
Comments and Suggestions
|last modified 07/09/2010 email Fermilab|
|Security, Privacy, Legal|